OpenStudy (anonymous):

Find the exact value of cos(195 degrees)

OpenStudy (anonymous):

I reckon your calculator could do that for you...

OpenStudy (anonymous):

i kn but we have to do it without a calculator

OpenStudy (mathstudent55):

\( \cos 195^o = \cos (150^o + 45^o)\) Use the angle addition identity.

OpenStudy (mathstudent55):

The calculator won't give an exact answer.

OpenStudy (kropot72):

You will find how to calculate the exact value at the following link: http://www.maths.surrey.ac.uk/hosted-sites/R.Knott/Fibonacci/simpleTrig.html#1575